Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 165     Entry time: 13 Sep 2024, 23:17
Author: Adam Round 
Group: SPB 
Subject: Shift end 

On Shift: TS, FT, CK, AR, PS

X-ray delivery
8.951 keV, 1.6 mJ up to 200 pulses at 0.5 MHz (intesity drops off at end of train)

Achievements/Observations
Beam trajectory and focusing re-established at 8.951 keV
Temporary beamstop verified position for sweeping - confirmed by eric
Beam sweeping in synchronisation with pulse trains observed!
slit window assembly remounted and beam is going through to back of hutch in perparation for optimisation next shft

ELOG V3.1.4-7c3fd00